Scenario:

Your 21-year old niece, who is graduating from college next month, asked for your advice – should she:

Invest $5,000 per year into her retirement fund (9 payments) from age 22 – 30?

Or wait until she turns 31 and invest $5,000 every year through age 65 (35 payments)?

Share your niece's age 65 balance from part A. for both scenarios with the class. What did you find? Share your findings from part B. What did you find? Discuss the implications for her goal of becoming financially independent.
Answer #1

The Question focuses on Study of Two Scenarios viz.,

1. Invest $5,000 per year into her retirement fund (9 payments) from age 22 – 30

Or

2. Wait until she turns 31 and invest $5,000 every year through age 65 (35 payments)

This should be decided based on the Present value of the Cash Outflow at different point of time and reaching at a conclusion Based on Net Cash Outflow.

Present value is the current worth of a future sum of money or stream of cash flow given a specified rate of return.

Future cash flows are discounted at the discount rate.

Determining the appropriate discount rate is the key to properly valuing future cash flows, whether they are earnings or obligations.Let us Assume the discount rate as 10% for Simplified Calculation purpose and for Study of Given two Scenarios.

1. Invest $5,000 per year into her retirement fund (9 payments) from age 22 – 30 :

Since in the absence of information related to timing of investment, it is assumed that it is done in the beginning of the year.

The Calculation as per Discounted Cash Outflow is presented as follows:

Age Investment(Cash Outflow)($)(i) PVF @ 10%(ii) (i)*(ii)
22 5000 0.909 4545.45
23 5000 0.826 4132.23
24 5000 0.751 3756.57
25 5000 0.683 3415.07
26 5000 0.621 3104.61
27 5000 0.564 2822.37
28 5000 0.513 2565.79
29 5000 0.467 2332.54
30 5000 0.424 2120.49
Total 45000 5.759 28795.12

PVF = Present Value factor at rate of 10%. formula for calculating PVF = 1/(1+r)n Where, r = discount rate and n = year

For first year 22, PVAF @ 10% = 1/(1.1)1 = 0.909

For Second Year 23, PVAF @ 10% = 1/(1.1)2 = 0.826

The Total Cash outflow discounted as on today(at the time, when Cousin age is 21 years) in above case is $28,795.12

The Above Amount of Present cash flows can also be calculated by using Cumulative Factor of Present value from year 22 to 30 as follows (This calculation can only done when the amount of investment p.a. is constant):

Present value of cash outflow from 22 to 30 years as on today = $5000 * 5.759 (As calculated above in the total of PVAF)

= $28,759.12

2. Wait until she turns 31 and invest $5,000 every year through age 65 (35 payments) :

Now, in this scenario, we can directly calculate Present value of future cash flows from year 31 to 65 (total 35 payments) ny using Cumulative Present Value factor from tenth Year from now(i.e. 31 years minus 21 years) upto Fourty-Fourth year(i.e. 65 year minus 21 years) which comes to 4.090 (Refer Working Note at the end of this answer)

Present value of cash outflow from 31 to 45 years as on today = $5000 * 4.090

= $20,450

Conclusion:

Since the Present value of cash flow is less in scenario 2 by $ 8,309 (i.e. 28,759-20,450), It is advised to choose for Scenario 2 i.e. wait until she turns 31 and invest $5,000 every year through age 65 (35 payments)

The implications for her goal of becoming financially independent:

Since the Present value of cash flow is less in scenario 2 by $ 8,309 as calculated above.There is a huge difference in amount invested.in first scenario, the total cash outflow in terms of $5000 per annum of 9 payments is calculated as 45000 while it is 175000 (5000*35 payments). Therefore, as becoming financially independent criteria as concerned, Choosing Scenario 2 i.e. to wait until she turns 31 and invest $5,000 every year through age 65 (35 payments) is more beneficial as more funds are invested which Directly results into increase in returns.

Working note:

Age Investment(i) PVAF @ 10%(ii) (i)*(ii)
31 5000 0.385543 1927.716
32 5000 0.350494 1752.469
33 5000 0.318631 1593.154
34 5000 0.289664 1448.322
35 5000 0.263331 1316.656
36 5000 0.239392 1196.96
37 5000 0.217629 1088.146
38 5000 0.197845 989.2233
39 5000 0.179859 899.2939
40 5000 0.163508 817.54
41 5000 0.148644 743.2181
42 5000 0.135131 675.6529
43 5000 0.122846 614.2299
44 5000 0.111678 558.3908
45 5000 0.101526 507.628
46 5000 0.092296 461.48
47 5000 0.083905 419.5273
48 5000 0.076278 381.3884
49 5000 0.069343 346.7167
50 5000 0.063039 315.197
51 5000 0.057309 286.5428
52 5000 0.052099 260.4934
53 5000 0.047362 236.8122
54 5000 0.043057 215.2838
55 5000 0.039143 195.7126
56 5000 0.035584 177.9205
57 5000 0.032349 161.7459
58 5000 0.029408 147.0417
59 5000 0.026735 133.6743
60 5000 0.024304 121.5221
61 5000 0.022095 110.4746
62 5000 0.020086 100.4315
63 5000 0.01826 91.30136
64 5000 0.0166 83.00123
65 5000 0.015091 75.45567
Total 175000 4.090065 20450.32